OpenStudy (anonymous):

How many moles of a gas sample are in 2.5 L container at 373 K and 203 kPa? 1.11 moles 0.66 moles 0.33 moles 0.17 moles

OpenStudy (jfraser):

The Ideal Gas Law is most useful here:\[P*V = n * R * T\] The value of R is\[R = 0.08205 \frac{L*atm}{mol*K}\] Rearrange to solve for n\[n = \frac{P*V}{R*T}\]Since the units of R have atm in them, you need to convert the kPa to atm.\[230kPa * (\frac{1atm}{101.3kPa})\]plug and chug.